OpenStudy (anonymous):

What is 5 x 1 3/5? I got 8, is that correct?

OpenStudy (welshfella):

write 1 3/5 as an improper fraction = 8/5 5 * 8/5 = 8

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

ok thanks

OpenStudy (welshfella):

welcome